OpenStudy (jaylelile):

give the slope-intercept form for the points (-4, 7) and (1,-3).

OpenStudy (anonymous):

-3-7/1+4=-10/5=-2

OpenStudy (anonymous):

the slope formula is y2-y1/x2-x1

OpenStudy (jaylelile):

yes I know that the slope is -5 over -10. But I don't know how to write it in slope intercept form (y=mx+b)....

OpenStudy (anonymous):

ok give me a minute

OpenStudy (anonymous):

excuse me i can't help you((

OpenStudy (triciaal):

so you know how to find the slope correct?

OpenStudy (jaylelile):

yes my slope is -5 over -10

OpenStudy (triciaal):

2 approaches one approach the slope is the same for any 2 points on the line so using one of the given points and (x, y) for any point equate to the same slope and simplify. rearrange to put in the form y = mx + b alternative approach using the calculated slope and one of the given points substitute to calculate b then write the equation for any point (x, y) with the calculated m and b

OpenStudy (triciaal):